maksolom7:

Triangle 2004-06-02-01-00_files/i0310000.jpg has vertices of A(–6, 7), B(4, –1), and C(–2, –9). Find the length of the median from 2004-06-02-01-00_files/i0310001.jpg in triangle 2004-06-02-01-00_files/i0310002.jpg.

Eiwoh2:

Sorry, we cannot see the image you tried to show us.

dude:

You can use the "Attach File" feature to add images

ratnakermehta20:

I hope the equations (eqns.) of the medians are required. I will work out the equation of the median AD through the vertex A(-6,7), where, D is the mid-point of the side BC. We have, B=B(4,-1) and C(-2,-9). Clearly, D=D((4-2)/2,(-1-9)/2)=D(1,-5). Having A(-6,7) and D(1,-5), the slope of AD is {7-(-5)}/(-6-1)=-12/7. Using the slope-point form eqn. of line, we get, the following eqn. of the median AD : y-7=-12/7(x-(-6)), i.e., 7y-49=-12x-72, or, 12x+7y+23=0. The other eqns. of the medians through the vertices B and C can be found similarly. Enjoy Maths.!